A 105-kg rubber mat is pushed on a dry concrete floor. What is the coefficient of static friction, if the force required to start the mat sliding is 1030 N?


1
Expert's answer
2021-03-14T19:14:01-0400

Applying the Newton’s Second Law of Motion, we get (at the moment when the mat begins to slide the acceleration is zero):


"F_{app}-F_{s.fr.}=0,""F_{appl}=F_{s.fr.}=\mu_sN=\mu_s mg,""\mu_s=\dfrac{F_{app}}{mg}=\dfrac{1030\ N}{105\ kg\cdot9.8\ \dfrac{m}{s^2}}=1.0"